OpenStudy (anonymous):

Help!! A store displays six computers on a shelf side–by–side. Find the total length of the computers on the table if the first computer was eight inches wide and the width of each successive computer is four inches longer than the previous one. Write in Sigma Notation.

OpenStudy (ash2326):

first computer 's length L1=8 second's length L2=8+4=8+4(2-1) so length of nth computer is Ln=8+4(n-1) so total length is L \[\sum_{1}^{8} (8+4(n-1))\]

OpenStudy (ash2326):

sorry i thought there are 8 computers , n should be from 1 to 6 \[L=\sum_{1}^{6}(8+4(n-1))\]

OpenStudy (anonymous):

Thank you soo much!!